# System 0/1/2/3: Quad-process theory for multi-timescale embodied collective cognitive systems

### v2

@@ 5.3 Collective World Models and Robotics Foundation Models
Recent advances in robotics have highlighted the development of robotics foundation models and vision-language-action (VLA) models, i.e., language-conditioned end-to-end robot controllers [2, 73, 23, 162, 70]. Although these approaches build upon traditional methods of policy learning through neural networks and position-based control, they increasingly emphasize the significance of integrating the influences of Systems 0 and 3. System 0 introduces the importance of passive dynamics in robotics, while System 3 focuses on how language, shaped by social interactions, plays a critical role in robot control. These developments position robotics foundation models as a unifying research domain that bridges System 0/1/2/3. VLA models exemplify how language”s socially constructed properties can guide robotic behavior.
In CPC, an agent’s physical form significantly influences its language structure. Consequently, robots—with physical forms distinct from humans—might develop unique linguistic structures tailored to their specific embodiments. This highlights the need for EmLang models that account for symbol and language emergence in the context of robotics [138]. Research in this area is becoming increasingly vital, as it seeks to integrate physical embodiment, social interaction, and cognitive dynamics into cohesive frameworks.
Additionally, System 3 emphasizes that language is not merely a medium for representing world models but serves as a collective world model itself [141]. Language encapsulates the shared experiences and knowledge of communities, functioning as a unifying representation of collective understanding. By adopting the perspective of System 0/1/2/3, these discussions gain clarity and provide a more comprehensive foundation for advancing robotics and AI. This integration opens pathways for innovative research on collective world models and their application in robotics.

### v3

@@ 5.3 Collective World Models and Robotics Foundation Models
Recent advances in robotics have highlighted the development of robotics foundation models and vision-language-action (VLA) models, i.e., language-conditioned end-to-end robot controllers (arai2024covla; kim2024openvla; dey2024revla; zhen20243d; Kawaharazuka16092024). Explicit integration of Systems 1 and 2 into robotics foundation models and VLAs can be found in recent studies (zawalski2024robotic; chen2025fast; lin2025onetwovla; intelligence2025pi). Although these approaches build upon traditional methods of policy learning through neural networks and position-based control, they are gradually realizing the significance of integrating the influences of Systems 0 and 3. System 0 introduces the importance of passive dynamics in robotics, while System 3 focuses on how language, shaped by social interactions, plays a critical role in robot control. These developments position robotics foundation models as a unifying research domain that bridges System 0/1/2/3. VLA models exemplify how language”s socially constructed properties can guide robotic behavior.
In CPC, an agent’s physical form significantly influences its language structure. Consequently, robots—with physical forms distinct from humans—might develop unique linguistic structures tailored to their specific embodiments. This highlights the need for EmLang models that account for symbol and language emergence in the context of robotics (taniguchi2016symbol). Research in this area is becoming increasingly vital, as it seeks to integrate physical embodiment, social interaction, and cognitive dynamics into cohesive frameworks.
Additionally, System 3 emphasizes that language is not merely a medium for representing world models but serves as a collective world model itself (taniguchi2024generative). Language encapsulates the shared experiences and knowledge of communities, functioning as a unifying representation of collective understanding. By adopting the perspective of System 0/1/2/3, these discussions gain clarity and provide a more comprehensive foundation for advancing robotics and AI. This integration opens pathways for innovative research on collective world models and their application in robotics.